# Event Camera Meets Resource-Aware Mobile Computing:

### v3

@@ Abstract.
With the increasing complexity of mobile device applications, these devices are evolving toward high agility. This shift imposes new demands on mobile sensing, particularly in achieving high-accuracy and low-latency. Event-based vision has emerged as a disruptive paradigm, offering high temporal resolution and low latency, making it well-suited for high-accuracy and low-latency sensing tasks on high-agility platforms. However, the presence of substantial noisy events, lack of stable, persistent semantic information, and large data volume pose challenges for event-based data processing on resource-constrained mobile devices. This paper surveys the literature from 2014 to 2025 and presents a comprehensive overview of event-based mobile sensing, encompassing its fundamental principles, event abstraction methods, algorithm advancements, and both hardware and software acceleration strategies. We discuss key applications of event cameras in mobile sensing, including visual odometry, object tracking, optical flow, and 3D reconstruction, while highlighting challenges associated with event data processing, sensor fusion, and real-time deployment. Furthermore, we outline future research directions, such as improving the event camera with advanced optics, leveraging neuromorphic computing for efficient processing, and integrating bio-inspired algorithms. To support ongoing research, we provide an open-source Online Sheet with recent developments. We hope this survey serves as a reference, facilitating the adoption of event-based vision across diverse applications.
Mobile Sensing; Event Camera; Event-based Vision
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computer systems organization Real-time systemsccs: Computer systems organization Embedded systemsccs: Computer systems organization Robotics
@@ 1. Introduction
Figure 1. Mobile devices are used in various applications, with key tasks including state estimation, environment perception, and understanding device-environment interactions. As devices become more agile, mobile sensing faces higher demands for accuracy and latency. This requires tight coordination between sensors and algorithms: (i) sensors must capture high-precision data with minimal delay; (ii) algorithms must efficiently process data within resource constraints. Traditional sensors fall short of these needs, whereas event cameras—capable of asynchronously capturing pixel-level intensity changes with microsecond latency—offer transformative potential. This survey presents a comprehensive review of event cameras and the development of efficient, accurate algorithms.
@@
Mobile sensing. With ongoing advancements in sensor technology and the proliferation of sophisticated computing capabilities within embedded systems, mobile devices (e.g., drones and autonomous vehicles) have emerged as the most groundbreaking innovations in recent years [1, 2, 3, 4, 5]. As illustrated in Fig.1, these devices are increasingly deployed in a variety of novel applications, including last-mile delivery [5], industrial inspection [6, 7, 8], rapid relief-and-rescue [9], aerial imaging [10] and sky networking [11], particularly within smart city scenarios. To perform these tasks, which require extensive interaction with the external environment, mobile devices must possess the ability to: $(i)$ awareness their own state, including location and orientation [12, 13, 14], $(ii)$ comprehend their surroundings, (e.g., environmental structure and map) [15], and $(iii)$ understand their relationship with the environment, such as the spatio-temporal relationships between mobile devices and objects within their surroundings [16]. Achieving these capabilities has become a focal point of interest within the mobile computing community.
@@
High agility trend of mobile devices. In context of smart city scenarios, the demand for mobile devices is progressively increasing, with their mission profiles evolving toward execution of 4D tasks, which are characterized as deep, dull, dangerous, and dirty [17]. The expanding scale of cities also necessitates that mobile devices complete various tasks within shorter time frames, driving their evolution toward high-speed operation [1, 18]. Consequently, the development of mobile devices is exhibiting new trends toward high agility. For instance, DJI’s industrial inspection drones can cruise at speeds of 21 $m/s$ [19], while Wing’s delivery drones fly at 30 $m/s$ to deliver packages [20].
@@
New challenges for mobile sensing. As mobile devices evolve toward high agility design, mobile sensing is required to advance toward high accuracy and low latency, enabling these mobile devices to perceive their state and surroundings in millimetermillimeter-level accuracy with millisecondmillisecond-level latency, thereby facilitating faster responses and more precise adjustments. This evolution establishes new objectives for the sensors and data processing algorithms involved in mobile sensing tasks: $(i)$ On the sensor input front, it is essential to acquire higher accuracy raw data with lower latency. $(ii)$ On data processing algorithms front, efficient processing of sensor measurements is essential to enhance accuracy in mobile sensing tasks while optimizing performance on resource-constrained platforms.
@@
Existing sensors for mobile sensing. However, existing sensors increasingly struggle to meet the high accuracy and low latency demands of mobile sensing, especially for environmental perception and interaction in high-agility devices. (i) Radar-based solutions employ sensors such as LiDAR [1, 21], mmWave radar [22], and ultrasound radar, which emit signals and estimate distances based on reflections. These methods track distance changes to update device positions and infer spatial relationships. However, they suffer from either high latency or limited accuracy: LiDAR achieves millimeter-level accuracy but requires point accumulation into frames at low frequencies (e.g., 10 Hz), introducing delays up to 100 ms; mmWave radar offers millisecond latency but lacks sufficient spatial resolution for millimeter accuracy. (ii) Camera-based methods use monocular [23, 24] and stereo cameras [25] for self-localization and environment mapping via SLAM [26]. Yet, these approaches are computationally demanding and limited by low temporal resolution (<30 Hz), relatively high latency (>30 ms), and standard dynamic range (60 dB), making them inadequate for high-agility mobile platforms.
@@
[27]
[28]
[29]
[30]
[31]
@@
[…]
Event-based data processing algorithms. Event cameras offer high-accuracy, low-latency data acquisition but face three main processing challenges: (i) Event cameras’ sensitivity to illumination causing significant noise, (ii) Event data lacks stable, persistent texture and cannot provide consistent information, as events are generated only at image edges depending on motion and scene texture, complicating feature extraction and long-term data association. (iii) Large data volume leading to high computational load on mobile devices. Efficient and accurate event data processing is thus crucial for resource-limited, agile mobile platforms to perceive their state and environment quickly. This survey reviews event processing algorithms across six key stages: event representation, denoising, filtering and feature extraction, matching, mapping, and hardware/software acceleration.
Difference between existing surveys. As shown in Fig.2, this survey extends previous surveys by focusing on how event cameras enable high-agility, resource-constrained mobile devices to achieve high-accuracy and low-latency self-state estimation and environmental understanding and extending prior surveys by incorporating literature from 2023–2025. It outlines the processing workflow of event data and reviews the advancements at each stage of this workflow. Using the key metrics of accuracy and efficiency in mobile computing, we summarize various methods in each stage to provide a deeper analysis of cutting-edge research, and assess each stage under latency, accuracy, and power constraints. It also covers work on event-based hardware and software acceleration, offering insights for deploying event cameras on resource-constrained mobile devices. Finally, we discuss applications of event cameras on mobile platforms.
@@
Contribution. The main contribution of the survey paper is summarized as follows.
$(1)$ We extend prior surveys by presenting a comprehensive review of how event cameras empower high-agility, resource-constrained mobile devices to achieve accurate, low-latency self-state estimation and environmental sensing.
$(3)$ We highlight unique advantages of event cameras in mobile sensing, as well as the specific challenges they encounter.
$(4)$ We categorize event stream processing methods into distinct stages and present a comprehensive review of each, covering event stream representation, data processing algorithms, acceleration strategies, and applications on event-based mobile platforms. Furthermore, we evaluate these stages with respect to latency, accuracy, and power constraints. $(5)$ We present our insights and solutions for future trends, with a particular focus on bio-inspired event camera hardware design, algorithm development, and hardware-software co-optimization techniques.
Online resource. This survey presents a comprehensive review of event-based sensing systems, focusing on key technological advancements and practical applications. To further support the research community, we have established an open-source Online Sheet  [^1] Event-based mobile sensing resource, which is adapted from [38]. This online sheet will be regularly updated, ensuring access to the latest developments and fostering continued innovation in event-based sensing systems.
Organization. Fig.2 illustrates the survey structure. Sec. 2 introduces event camera fundamentals, including principles, hardware, products, datasets, and their pros and cons on mobile devices. Sec. 3 covers event stream abstraction methods, while Sec. 4 reviews event processing algorithms such as denoising, filtering, matching, and mapping. Sec. 5 discusses hardware/software acceleration for resource-constrained devices, and Sec. 6 highlights mobile applications of event cameras. Sec. 7 outlines future directions, with conclusions in Sec. 8.
@@ 2.1. Event generation model
Unlike conventional cameras that capture images at fixed intervals, event cameras operate asynchronously by detecting changes in log intensity at individual pixels, generating events only when significant changes occur. This characteristic enables event cameras to achieve exceptionally high temporal resolution and effectively mitigate motion blur, particularly in scenarios involving fast-moving objects.
@@
Each event is defined by the pixel location where the change occurs, the timestamp, and the polarity. Formally, an event can be represented as $e_{k}=(x_{k},t_{k},p_{k})$, where $e_{k}$ denotes the event, $x_{k}$ specifies the pixel location, $t_{k}$ represents the timestamp, and $p_{k}$ indicates the polarity of the change. The change in log intensity is given by $\Delta L(x_{k},t_{k})=L(x_{k},t_{k})-L(x_{k},t_{k}-\Delta t_{k})$, where $L(x_{k},t_{k})$ represents the log intensity at pixel $x_{k}$ and time $t_{k}$. As shown in Fig.3, an event is triggered only when $|\Delta L(x_{k},t_{k})|$ exceeds a predefined threshold $C$. The polarity $p_{k}$ is assigned as +1+1 if $\Delta L(x_{k},t_{k})>0$, and 1-1 otherwise. This event-driven paradigm substantially reduces redundant data, enhancing processing efficiency, conserving computational resources, and enabling deployment in resource-constrained systems such as embedded devices [39]. In practice, threshold $C$ can be adjusted to meet specific application needs and impacts event camera performance. A high $C$ reduces sensitivity, may missing subtle changes, while a low $C$ increases noise-triggered events, causing redundancy.
@@ 2.2. Hardware design
DVS event camera [40]. The Dynamic Vision Sensor (DVS), inspired by the silicon retina, detects brightness changes via capacitance coupling and resets after each measurement. It outputs only changes, enabling smaller pixel sizes but restricting output to event data, which limits information extraction in static scenes.
ATIS event camera [41]. The Active Time-Image Sensor (ATIS) uses subpixels to measure absolute brightness, doubling pixel area compared to DVS but enabling wide dynamic and static ranges for robust imaging under extreme lighting. Its limitation lies in potential misalignment between absolute brightness (averaged across pixels) and event data (triggered per pixel), especially during high-speed motion.
@@
Indicator Original event camera[42] DVS [40] ATIS [41] DAVIS [43]
@@
DAVIS event camera [43]. The Dynamic and Active Vision Sensor (DAVIS) is capable of outputting both absolute brightness and event-based data. In DAVIS, pixels and subpixels share the same sensor, enabling a more compact design. As a result, the pixel area is smaller than that of the ATIS, with only a modest 5% increase in size compared to the DVS. However, the sampling speed of the DAVIS circuit is slower than that of the DVS circuit.
Beyond foundational architectures, modern event cameras are optimized for mobile platforms. Key trends include more versatile readouts (e.g., hardware-generated event-accumulation views in the CeleX-V [44]), mobile-friendly interfaces like MIPI CSI-2 (on the CeleX5-MIPI [45]), and ultra-low power optimization specifically targeting battery-powered mobile devices (e.g., Prophesee’s GENX320-class sensors [46]). A detailed comparison of these commercial products is presented in Section 2.5.
2.3. Advantages of event cameras in mobile sensing
@@
High temporal resolution. The high agility of mobile devices induces rapid environmental changes, often causing motion blur in frame cameras and limiting the responsiveness of radar- or camera-based solutions due to their low spatio-temporal resolution. Temporal resolution, defined as the interval between consecutive samples, improves with shorter intervals. Event cameras, with microsecond-level resolution and motion-blur-free sensing, enable timely detection of both environmental dynamics and device state.
@@
Low sensing latency. The high agility of mobile devices demands rapid awareness of both environmental and self-state changes. Sensing latency, defined as the time required for a sensor to respond to environmental variations, is relatively long in frame cameras due to their global exposure time (20 ms), which delays reactions and increases collision risk. In contrast, event cameras employ independent pixels that trigger events immediately upon brightness changes, achieving sub-millisecond latency and enabling mobile devices to detect and respond to changes instantaneously.
@@
High dynamic range (HDR). Mobile devices are increasingly used in challenging environments, such as low-light nighttime and bright daytime settings, requiring reliable sensing across varying lighting conditions. Standard frame cameras have a dynamic range of about 60 $dB$, making them less effective in extreme lighting. Event cameras operate on a logarithmic scale with independent pixels, offering a very high dynamic range with >120 $dB$. This enables them to adapt to both extremely dark and bright conditions, making mobile devices suitable for a wider range of scenarios.
@@
Low power consumption. Mobile devices in complex urban environments often prioritize efficiency but face limited computation and power. Frame cameras require heavy processing and energy, whereas event cameras transmit only brightness changes, reducing redundant data and easing computational and power demands to improve efficiency.
2.4. Challenges of event cameras in mobile sensing

### v4

# Event Camera Meets Mobile Embodied Perception:

@@ Abstract.
With the evolution of mobile embodied intelligence, agents such as drones and autonomous robots are transitioning toward high agility. This shift imposes stringent demands on embodied perception, requiring high-accuracy and low-latency feedback loops for reliable interaction. Event-based vision has emerged as a transformative paradigm. Its microsecond-level temporal resolution and high dynamic range render it ideal for embodied perception tasks on high-agility mobile platforms. However, asynchronous nature, substantial noise, lack of persistent semantic information, and large data volume pose challenges for processing on resource-constrained mobile agents. This paper surveys the literature from 2014-2025 and presents a comprehensive overview of event-based mobile embodied perception. We organize review around four key pillars: event abstraction methods, perception algorithm advancements, hardware and software acceleration strategies, and mobile applications. We discuss critical tasks including visual odometry, object tracking, optical flow, and 3D reconstruction, while highlighting challenges associated with sensor fusion and real-time deployment. Furthermore, we outline future research directions, such as improving event cameras with advanced optics and leveraging neuromorphic computing for efficient processing. To support ongoing research, we provide an open-source Online Sheet with recent developments. We hope this survey serves as a reference, facilitating adoption of event-based vision across diverse mobile embodied applications.
Embodied AI, Mobile Embodied Perception; Event Camera; Event-based Vision
This paper was supported by the Natural Science Foundation of China under Grant 62371269, Guangdong Innovative and Entrepreneurial Research Team Program 2021ZT09L197, Shenzhen Low-Altitude Airspace Strategic Program Portfolio Z253061 and Meituan Academy of Robotics Shenzhen.
copyright: acmcopyrightjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computer systems organization Real-time systemsccs: Computer systems organization Embedded systemsccs: Computer systems organization Roboticsccs: Computing methodologies Computer visionccs: Computing methodologies Bio-inspired approaches
@@ 1. Introduction
Figure 1. Mobile agents are used in various applications, with key tasks including state estimation, environment perception, and understanding agent-environment interactions. As agents become more agile, mobile perception faces higher demands for accuracy and latency. This requires tight coordination between sensors and algorithms: (i) sensors must capture high-precision data with minimal delay; (ii) algorithms must efficiently process data within resource constraints. Traditional sensors fall short of these needs, whereas event cameras, capable of asynchronously capturing pixel-level intensity changes with microsecond latency, offer transformative potential. This survey presents a comprehensive review of event cameras and development of efficient algorithms.
@@
In the era of embodied AI, mobile agents, ranging from drones and autonomous vehicles to humanoid robots, are expected to operate autonomously in the wild [chen2024ddl, zhou2022swarm, jian2023path]. Unlike passive perception systems, these embodied agents must actively perceive their environment to make real-time decisions, forming a tight coupling between perception and action [ren2025safety, wang2025ultra]. As illustrated in Fig.1, these agents are increasingly deployed in a variety of applications, including last-mile delivery [wang2025ultra], industrial inspection [wang2024transformloc, xu2022swarmmap, liu2024mobiair], rapid relief-and-rescue [chen2024soscheduler], aerial imaging [leng2024recent] and sky networking [xu2024scalable], particularly within smart city scenarios. To perform these tasks, which require extensive interaction with the external environment, mobile agents must possess the ability to: $(i)$ perceive their own state, including location and orientation [wang2025aerial, chen2023self, wang2022h], $(ii)$ comprehend their surroundings, (e.g., environmental structure and map) [cui2024alphalidar, zhao2025flight], and $(iii)$ understand their relationship with the environment, such as the spatio-temporal relationships between mobile agents and objects within their surroundings [he2022automatch, wang2025enabling]. Achieving these capabilities has become a focal point of interest within mobile embodied intelligence community.
@@
In the context of smart city scenarios, the demand for mobile agents is progressively increasing, with their mission profiles evolving toward execution of 4D tasks, which are characterized as deep, dull, dangerous, and dirty [francis2022gas]. The expanding scale of cities also necessitates that mobile agents complete various tasks within shorter time frames, driving their evolution toward high-speed operation [ren2025safety, floreano2015science]. Consequently, the development of mobile agents is exhibiting new trends toward high agility. For instance, DJI’s industrial inspection drones can cruise at speeds of 21 $m/s$ [DJIMAVIC4], while Wing’s delivery drones fly at 30 $m/s$ to deliver packages [Wing].
@@
As mobile agents evolve toward high agility design, mobile embodied perception is required to advance toward high accuracy and low latency, enabling these mobile agents to perceive their state and surroundings in millimetermillimeter-level accuracy with millisecondmillisecond-level latency, thereby facilitating faster responses and more precise adjustments. This evolution establishes new objectives for the sensors and data processing algorithms involved in mobile embodied perception tasks: $(i)$ On the sensor input front, it is essential to acquire higher accuracy raw data with lower latency. $(ii)$ On the data processing algorithms front, efficient processing of sensor measurements is essential to enhance accuracy in mobile perception tasks while optimizing performance on resource-constrained mobile platforms.
@@
However, existing sensors increasingly struggle to meet the high accuracy and low latency demands of mobile perception, especially for environmental perception and interaction in high-agility agents. (i) Radar-based solutions employ sensors such as LiDAR [ren2025safety, zhu2024swarm], mmWave radar [sie2024radarize], and ultrasound radar, which emit signals and estimate distances based on reflections. These methods track distance changes to update agent positions and infer spatial relationships. However, they suffer from either high latency or limited accuracy: LiDAR achieves millimeter-level accuracy but requires point accumulation into frames at low frequencies (e.g., 10 Hz), introducing delays up to 100 ms; mmWave radar offers millisecond latency but lacks sufficient spatial resolution for millimeter accuracy. (ii) Frame camera-based methods use monocular [zhang2022edgeis, cao2023edge] and stereo cameras [qin2019a] for self-localization and environment mapping via SLAM [li2024tmc_edgeslam2]. However, these approaches are computationally intensive and constrained by low temporal resolution (<30 Hz), high latency (>30 ms), motion blur, and limited dynamic range (60 dB). These limitations decouple perception from the continuous dynamics of the real world, rendering them unsuitable for high-agility mobile agents. Embodied perception demands a sensor that is natively responsive to motion.
[…]
Event cameras offer high-accuracy, low-latency data acquisition but face three main processing challenges: (i) Event cameras’ sensitivity to illumination causing significant noise, (ii) Event data lacks stable, persistent texture and cannot provide consistent information, as events are generated only at image edges depending on motion and scene texture, complicating feature extraction and long-term data association. (iii) Large data volume leading to high computational load on mobile agents. Efficient and accurate event data processing is thus crucial for resource-constrained, agile mobile agents to perceive their state and environment. This survey reviews event processing algorithms across six stages: event representation, denoising, filtering and feature extraction, matching, mapping, and hardware/software acceleration.
@@
[gallego2020event]
[zheng2023deep]
[chakravarthi2024recent]
[novo2024neuromorphic]
[shariff2024event]
@@
As shown in Tab. 1, this survey extends previous surveys by focusing on how event cameras enable high-agility, resource-constrained mobile agents to achieve high-accuracy and low-latency self-state estimation and environmental understanding and extending prior surveys by incorporating literature from 2023–2025. It outlines the processing workflow of event data and reviews the advancements at each stage of this workflow. Using the key metrics of accuracy and efficiency in mobile computing, we summarize various methods in each stage to provide a deeper analysis of cutting-edge research, and assess each stage under latency, accuracy, and power constraints. It also covers work on event-based hardware and software acceleration, offering insights for deploying event cameras on resource-constrained mobile agents. Finally, we discuss applications of event cameras on mobile agents.
The main contribution of the survey paper is summarized as follows.
$(1)$ We extend prior surveys by presenting a comprehensive review of how event cameras empower high-agility, resource-constrained mobile agents to achieve accurate, low-latency self-state estimation and environmental perception.
$(3)$ We highlight unique advantages and specific challenges of event cameras in mobile embodied perception.
$(4)$ We categorize event stream processing methods into distinct stages and present a comprehensive review of each, covering event stream representation, data processing algorithms, acceleration strategies, and applications on event-based mobile agents. Furthermore, we evaluate these stages with respect to latency, accuracy, and power constraints.
$(5)$ We present our insights and solutions for future trends, with a particular focus on bio-inspired event camera hardware design, algorithm development, and hardware-software co-optimization techniques.
@@
This survey presents a comprehensive review of event-based embodied perception systems, focusing on key technological advancements and practical applications. To further support the research community, we have established an open-source Online Sheet  [^1] Event-based mobile embodied perception resource, which is adapted from [Recent]. This online sheet will be regularly updated, ensuring access to latest developments and fostering continued innovation in event-based embodied perception systems.
Fig.2 illustrates the survey structure. Sec. 2 introduces event camera fundamentals, including principles, hardware, products, datasets, and their pros and cons on mobile agents. Sec. 3 covers event stream Abstraction methods, while Sec. 4 reviews event processing Algorithms such as denoising, filtering, matching, and mapping. Sec. 5 discusses hardware/software Acceleration for resource-constrained agents, and Sec. 6 highlights mobile Applications of event cameras. Sec. 7 outlines future directions, with conclusions in Sec. 8.
@@ 2.1. Event generation model
Unlike conventional frame cameras that capture images at fixed intervals, event cameras operate asynchronously by detecting changes in log intensity at individual pixels, generating events only when changes occur. This characteristic enables event cameras to achieve exceptionally high temporal resolution and effectively mitigate motion blur, particularly in scenarios involving fast-moving objects.
Each event is defined by the pixel location where the change occurs, the timestamp, and the polarity. Formally, an event can be represented as $e_{k}=(x_{k},t_{k},p_{k})$, where $e_{k}$ denotes the event, $x_{k}$ specifies the pixel location, $t_{k}$ represents the timestamp, and $p_{k}$ indicates the polarity of the change. The change in log intensity is given by $\Delta L(x_{k},t_{k})=L(x_{k},t_{k})-L(x_{k},t_{k}-\Delta t_{k})$, where $L(x_{k},t_{k})$ represents the log intensity at pixel $x_{k}$ and time $t_{k}$. As shown in Fig.3, an event is triggered only when $|\Delta L(x_{k},t_{k})|$ exceeds a predefined threshold $C$. The polarity $p_{k}$ is assigned as +1+1 if $\Delta L(x_{k},t_{k})>0$, and 1-1 otherwise. This event-driven paradigm substantially reduces redundant data, enhancing processing efficiency, conserving computational resources, and enabling deployment in resource-constrained systems such as embedded mobile agents [xu2023taming]. In practice, threshold $C$ can be adjusted to meet specific application needs and impacts event camera performance. A high $C$ reduces sensitivity, may missing subtle changes, while a low $C$ increases noise-triggered events, causing redundancy.
@@ 2.2. Hardware design
DVS event camera [lichtsteiner2008128]. The Dynamic Vision Sensor (DVS), inspired by the silicon retina, detects brightness changes via capacitance coupling and resets after each measurement. It outputs only changes, enabling smaller pixel sizes but restricting output to event data, which limits information extraction in static scenes.
ATIS event camera [posch2010qvga]. The Active Time-Image Sensor (ATIS) uses subpixels to measure absolute brightness, doubling pixel area compared to DVS but enabling wide dynamic and static ranges for robust imaging under extreme lighting. Its limitation lies in potential misalignment between absolute brightness (averaged across pixels) and event data (triggered per pixel), especially during high-speed motion.
DAVIS event camera [brandli2014240]. The Dynamic and Active Vision Sensor (DAVIS) is capable of outputting both absolute brightness and event-based data. In DAVIS, pixels and subpixels share the same sensor, enabling a more compact design. As a result, the pixel area is smaller than that of the ATIS, with only a modest 5% increase in size compared to the DVS. However, the sampling speed of the DAVIS circuit is slower than that of the DVS circuit.
Beyond foundational architectures, modern event cameras are optimized for mobile agents. Key trends include more versatile readouts (e.g., hardware-generated event-accumulation views in the CeleX-V [chen2019live]), mobile-friendly interfaces like MIPI CSI-2 (on the CeleX5-MIPI [celex5mipi_datasheet]), and ultra-low power optimization specifically targeting battery-powered mobile agents (e.g., Prophesee’s GENX320 [prophesee_genx320_datasheet]). A detailed comparison of commercial products is presented in Sec. 2.5.
@@
Indicator Original event camera[mahowald1994silicon] DVS [lichtsteiner2008128] ATIS [posch2010qvga] DAVIS [brandli2014240]
@@
2.3. Advantages of event cameras in mobile embodied perception
@@
High temporal resolution. The high agility of mobile agents induces rapid environmental changes, often causing motion blur in frame cameras and limiting the responsiveness of radar- or frame camera-based solutions due to their low spatio-temporal resolution. Temporal resolution, defined as the interval between consecutive samples, improves with shorter intervals. Event cameras, with microsecond-level resolution and motion-blur-free embodied perception, enable timely detection of both environmental dynamics and agent state.
@@
Low perception latency. The high agility of mobile agents demands rapid awareness of both environmental and self-state changes. Perception latency, defined as the time required for a sensor to respond to environmental variations, is relatively long in frame cameras due to their global exposure time (20 ms), which delays reactions and increases collision risk. In contrast, event cameras employ independent pixels that trigger events immediately upon brightness changes, achieving sub-millisecond latency and enabling mobile agents to detect and respond to changes instantaneously.
@@
High dynamic range (HDR). Mobile agents are increasingly used in challenging environments, such as low-light nighttime and bright daytime settings, requiring reliable embodied perception across varying lighting conditions. Frame cameras have a dynamic range of about 60 $dB$, making them less effective in extreme lighting. Event cameras operate on a logarithmic scale with independent pixels, offering a high dynamic range with >120 $dB$. This enables them to adapt to both extremely dark and bright conditions, making mobile agents suitable for a wider range of scenarios.
@@
Low power consumption. Mobile agents in complex urban environments often prioritize efficiency but face limited computation and power. Frame cameras require heavy processing and energy, whereas event cameras transmit only brightness changes, reducing redundant data and easing computational and power demands to improve efficiency.
2.4. Challenges of event cameras in mobile embodied perception